# WeFT: Weighted Entropy-driven Fine-Tuning for dLLMs

### v1

@@ 4.2 Main Results
As shown in Table 1, we compare the performance of WeFT and the standard SFT method across different training datasets and benchmarks. Compared with the base LLaDA-8B-Instruct model, both SFT and WeFT improve performance on all tasks. Notably, WeFT consistently achieves higher accuracy than SFT, with a 39% improvement on the s1K dataset, a 64% improvement on the s1K-1.1 dataset, and a 83% improvement on the openr1-3k dataset, demonstrating that our method effectively leverages token importance and structured information. These results indicate that WeFT provides robust and reliable performance gains across tasks and datasets, particularly for tasks requiring complex reasoning.
Model Sudoku Countdown GSM8K MATH500 Average
0-shot 0-shot 0-shot 0-shot
Base Model
LLaDA-8B-Instruct 5.5 16.0 76.7 32.4 32.6
s1K Dataset
+ SFT 4.6 (-0.9) 23.8 (+7.8) 78.8 (+2.1) 32.6 (+0.2) 34.9 (+2.3)
+ WeFT 7.8 (+2.3) 24.6 (+8.6) 78.0 (+1.3) 33.0 (+0.6) 35.8 (+3.2)
s1K-1.1 Dataset
+ SFT 4.2 (-1.3) 21.5 (+5.5) 78.4 (+1.7) 30.8 (-1.6) 33.7 (+1.1)
+ WeFT 6.8 (+1.3) 20.3 (+4.3) 78.1 (+1.4) 32.6 (+0.2) 34.4 (+1.8)
openr1-3k Dataset
+ SFT 6.3 (+0.8) 17.2 (+1.2) 77.3 (+0.6) 32.0 (-0.4) 33.2 (+0.6)
+ WeFT 6.9 (+1.4) 17.6 (+1.6) 77.5 (+0.8) 32.8 (+0.4) 33.7 (+1.1)
Table 1: Comparison of the performance of WeFT and SFT across different training datasets and benchmarks. The table shows that WeFT achieves substantial percentage improvements over SFT.
4.3 SFT Role in Subsequent RL
Reinforcement learning (RL) post-training has become a crucial stage in the training of large language models  (DeepSeek-AI, 2025). Supervised fine-tuning (SFT) cold-start plays an important role in preparing the model for subsequent RL, as SFT establishes a reliable initial policy which can then be further refined by RL objectives. Therefore, it is essential to verify whether models trained with WeFT can maintain their advantages during subsequent RL training. To this end, we fine-tune models on the s1K dataset using both WeFT and SFT, and then further train them with RL on the Sudoku dataset for comparison.
Figure 3: Reward curves of models cold-started with WeFT and SFT during subsequent reinforcement learning training. The curves are smoothed using a time-weighted EMA. As shown, the model initialized with WeFT achieves higher rewards.
As shown in Figure 3, models initialized with WeFT achieve higher rewards throughout RL training. Moreover, as reported in Table 2, the WeFT-trained model achieves significantly better evaluation results, with a 49% relative performance gain. These findings demonstrate that models cold-started with WeFTare highly effective in guiding subsequent RL learning.
Model LLaDA-8B-Instruct + SFT + WeFT + SFT + RL +WeFT + RL
Sudoku 5.5 4.6 7.8 8.9 13.3
Table 2: Evaluation results of models cold-started with WeFT and SFT during subsequent reinforcement learning training. As shown in the table, the model initialized with WeFT achieves significantly higher evaluation accuracy.
@@ 4.4 Ablation Study
In this subsection, we conduct ablation studies to validate the design choices of WeFT, focusing on three aspects: the necessity of using entropy as the metric, the validity of the theoretically derived loss function, and the empirical necessity of applying the square root of entropy during training.
4.4.1 the Necessity of Using Entropy as the Metric
To demonstrate the necessity of entropy-based weighting, we compared our method with a baseline using the negative log-likelihood (NLL). Unlike entropy-based weighting, NLL emphasizes tokens that appear less frequently rather than those that are difficult to predict or carry more information. Assuming the logits are $z_{i}$, the masking rate $\beta_{i}$ for the $i$-th token under the NLL setting is:
$\beta_{i}=-\log\bigl(\text{softmax}(z_{i})^{x_{0}^{i}}\bigr)$ (14)
@@
+ WeFT 7.8 (+2.3) 24.6 (+8.6) 78.0 (+1.3) 33.0 (+0.6) 35.8 (+3.2)
Table 3: Comparison between entropy-based weighting and NLL-based weighting. As shown in the table, the SFT algorithm with NLL-based weighting performs significantly worse than the entropy-based variant, and on many tasks even underperforms the original model.
@@
4.4.2 the Validity of the Theoretically Derived Loss Function
@@
$L=\sum_{i}\mathbb{E}_{t}\left[\mathbf{1}\!\left[x_{t}^{i}=\mathbf{M}\right]\frac{w_{i}}{t}\log\!\left(x_{0}^{i}\mid x_{t}\right)\right]$ (15)
@@
Table 4 presents a comparison between WeFT, WeFT (SW) (i.e., simple token-wise weighting using our square-root entropy as the weight), and Dream. Dream yields only minimal improvements and, in some cases, performs worse than standard SFT. WeFT (SW) provides some gains but still performs worse than SFT, whereas WeFT significantly outperforms SFT. This demonstrates that the diffusion loss derived from theoretical principles possesses superior properties.
@@
+ WeFT (SW) 5.3 (-0.2) 23.8 (+7.8) 75.9 (-0.8) 31.6 (-0.8) 34.2 (+1.6)
+ WeFT 7.8 (+2.3) 24.6 (+8.6) 78.0 (+1.3) 33.0 (+0.6) 35.8 (+3.2)
Table 4: Comparison between WeFT, WeFT (SW) (simple token-wise weighting using our square-root entropy as the weight), and Dream. Dream yields negligible improvements. WeFT (SW) provides some gains, but still performs worse than SFT, whereas WeFT significantly outperforms SFT.
4.4.3 the Empirical Necessity of Applying the Square Root of Entropy During Training
Below, we explain why we choose the square root of entropy rather than entropy as the metric. As shown in Table 5, we observed that directly using token weights based on raw entropy leads to large variations across tokens, which can significantly reduce training stability and even cause oscillations. To address this issue, we implement WeFT using the square root of entropy as the weighting factor. This modification effectively reduces the scale of the gradient norms, resulting in more stable training and improved overall convergence behavior.
@@ 4.5 Time Efficiency Analysis
Figure 4: Time efficiency analysis of SFT and WeFT.
Compared to SFT, WeFT introduces one additional forward pass per training step. To evaluate the computational overhead, we measured the time required to train both SFT and WeFT for 20 epochs on the s1K dataset using 4 H100 GPUs. As shown in Figure 4, the results show that WeFT incurs only a modest increase in training time (approximately 24% more than standard SFT), demonstrating that the performance gains of WeFT come with minimal additional computational cost.
@@ 5.1 Diffusion Large Language Models
Diffusion models, initially developed for continuous data, have recently been adapted to discrete sequences, including natural language. Early work such as Structured Denoising Diffusion Models (D3PMs) (Austin et al., 2021) extended discrete diffusion to text, showing competitive performance in character-level generation and large vocabulary language modeling. DiffusionBERT (He et al., 2023) combined masked language modeling with diffusion, improving generative quality over conventional masked LMs. Later, masked diffusion frameworks (Sahoo et al., 2024; Shi et al., 2024) introduced simplified and generalized training objectives that reduce complexity while maintaining performance, and support state-dependent masking strategies. Diffusion Large Language Models (dLLMs) including LLaDA (Nie et al., 2025) and Dream  (Ye et al., 2025) demonstrated that diffusion-based models can scale to billions of parameters, achieving strong context learning and instruction-following capabilities. Techniques such as variance-reduced preference optimization (Zhu et al., 2025) and reinforcement of lateral thought chains (Huang et al., 2025) further enhance generative stability and reasoning ability, while efficient perplexity bounds and ratio matching (Haxholli et al., 2025) improve sampling efficiency. More recently, works such as d1 (Zhao et al., 2025) and wd1 (Tang et al., 2025) have explored the possibility of applying reinforcement learning on top of existing dLLMs.
@@ 5.2 Weighted Methods in Language Model Learning
A complementary line of research focuses on improving large language models through weighted supervised fine-tuning (SFT) and reinforcement learning. Weighted SFT techniques aim to prioritize informative tokens. DFT (Wu et al., 2025) addresses the limited generalization of SFT by dynamically rescaling token-level objectives to stabilize gradient updates. Rho-1 (Lin et al., 2024) challenges the conventional uniform next-token prediction objective by selectively training on useful tokens identified via a reference model. Similarly, Luo et al. (Luo et al., 2023) propose a re-weighting strategy for active learning in NER, assigning dynamic smoothed weights to tokens to boost performance. In reinforcement learning, applying weighting to improve learning effectiveness is also a common strategy. (Cui et al., 2025) investigate the entropy dynamics of reinforcement learning for reasoning LLMs, and propose entropy-based weighting methods to prevent entropy collapse and encourage exploration. (Wang et al., 2025) highlight the critical role of high-entropy minority tokens in RLVR, demonstrating that restricting updates to these tokens significantly enhances reasoning capabilities. Together, these approaches demonstrate that token-level weighting is an important and effective way to improve the optimization of language models.
@@ 6 Conclusion
We introduced WeFT, a theoretically grounded weighted SFT algorithm for diffusion-based language models. By leveraging token-level entropy as a measure of importance, WeFT prioritizes high-uncertainty tokens and preserves the diffusion process properties. Experiments across multiple benchmarks show substantial and consistent improvements over standard SFT, with gains persisting into RL training, all at minimal computational overhead. These results highlight entropy-aware weighting as a simple yet powerful strategy for enhancing reasoning and stability in diffusion-based LLMs.
@@ Appendix A Proof of the Weighted SFT Loss
In order to prove this theorem, we will first prove two lemmas, and then proceed to prove the main theorem. Our proof follows that of (Ou et al., 2025), with the key difference that our $Q$ matrix incorporates varying masking rates $\beta$, whereas (Ou et al., 2025) assumes all masking rates are identical.
@@ A.1 Proof of Lemmas
$\frac{p_{t}(x_{t}^{1},\dots,{x^{\prime}}^{i}_{t},\dots,x_{t}^{d})}{p_{t}(x_{t}^{1},\dots,x^{i}_{t},\dots,x_{t}^{d})}=\frac{\exp(-\beta_{x_{t}^{i}}\,\bar{f}(t))}{1-\exp(-\beta_{x_{t}^{i}}\,\bar{f}(t))}\;p_{0}({x^{\prime}}_{t}^{i}\mid x_{t}^{\mathrm{unmasked}}).$ (18)
@@
$\frac{p_{t}(x_{t}^{1},\dots,{x^{\prime}}^{i}_{t},\dots,x_{t}^{d})}{p_{t}(x_{t}^{1},\dots,x^{i}_{t},\dots,x_{t}^{d})}=0.$ (19)
@@
$=$ $\prod_{i=0}^{n-1}(1+Q_{s_{i}}(x_{s_{i}},x_{s_{i}})(s_{i+1}-s_{i})+o(s_{i+1}-s_{i}))$
$=$ $\exp(\sum_{i=0}^{n-1}\ln(1+Q_{s_{i}}(x_{s_{i}},x_{s_{i}})(s_{i+1}-s_{i})+o(s_{i+1}-s_{i})))$
$=$ $\exp(\sum_{i=0}^{n-1}[f(s_{i})Q(x_{s_{i}},x_{s_{i}})(s_{i+1}-s_{i})+o(s_{i+1}-s_{i})])$
Now we select the the appropriate $n$ and $s_{1},\dots,s_{n-1}$ such that $n\rightarrow+\infty$ and $\max(s_{i+1}-s_{i})\rightarrow 0$, recalling the definition of integral, we obtain:
$\exp(\sum_{i=0}^{n-1}[f(s_{i})Q(x_{s_{i}},x_{s_{i}})(s_{i+1}-s_{i})+o(s_{i+1}-s_{i})])\rightarrow\exp(-\beta_{x_{0}}\bar{f}(t))$
@@
$p_{t|0}(x_{t}|x_{0})=\left\{\begin{array}[]{ll}\exp(-\beta_{x_{0}}\bar{f}(t))&\text{if }x_{t}=x_{0}\\ 1-\exp(-\beta_{x_{0}}\bar{f}(t))&\text{if }x_{t}=[M]\\ 0&\text{otherwise}\end{array}\right$.
Now, based on the above result, recall the independence of different dimensions in the diffusion process, we can get a stronger result:
@@
Suppose we have $x_{t}$ in which the tokens $a_{1},a_{2},...,a_{k}$ are masked and $b_{1},b_{2},...,b_{n-k}$ are unmasked, then:
$p_{t}(x_{t})=\prod_{i=1}^{k}(1-\exp(-\beta_{a_{i}}\bar{f}(t)))\prod_{j=1}^{n-k}\exp(-\beta_{b_{j}}\bar{f}(t))p_{0}(x_{t}^{\text{unmasked}})$ (20)
@@
$p_{t}([M][M]\dots[M]x_{t}^{k+1}x_{t}^{k+2}\dots x_{t}^{n})$
$=$ $\sum_{x_{0}}p_{t|0}([M][M]\dots[M]x_{t}^{k+1}x_{t}^{k+2}\dots x_{t}^{n}|x_{0})p_{0}(x_{0})$
$=$ $\sum_{x_{0}^{1}\dots x_{0}^{n}}p_{t|0}([M][M]\dots[M]x_{t}^{k+1}x_{t}^{k+2}\dots x_{t}^{n}|x_{0}^{1},x_{0}^{2},\dots x_{0}^{n})p_{0}(x_{0}^{1},\dots x_{0}^{n})$
$=$ $\sum_{x_{0}^{1}\dots x_{0}^{n}}\prod_{i=1}^{k}p^{k}_{t|0}([M]|x_{0}^{k})\prod_{i=k+1}^{n}p_{t|0}(x_{t}^{k}|x_{0}^{k})p_{0}(x_{0}^{1},\dots x_{0}^{n})$
$=$ $\sum_{x_{0}^{1}\dots x_{0}^{n}}\prod_{i=1}^{k}(1-\exp(-\beta_{i}\bar{f}(t)))\prod_{i=k+1}^{n}\exp(-\beta_{i}\bar{f}(t))p_{0}(x_{0}^{1},\dots x_{0}^{n})$
$=$ $\sum_{x_{0}^{k+1}\dots x_{0}^{n}}\prod_{i=1}^{k}(1-\exp(-\beta_{i}\bar{f}(t)))\prod_{i=k+1}^{n}\exp(-\beta_{i}\bar{f}(t))p_{0}(x_{0}^{k+1},\dots x_{0}^{n})$
Since $p_{0}(x_{0}^{k+1},\dots,x_{0}^{n})$ can be expressed as $p_{0}(x_{t}^{\text{unmasked}})$, the desired result follows.

### v2

@@ 4.2 Main Results
As shown in Table 1, we compare the performance of GIFT and the standard SFT method across different training datasets and benchmarks. Due to computational constraints, each of our experiments is the result of a single run. Notably, GIFT consistently achieves higher average accuracy than SFT, demonstrating that our method achieves effective training by leveraging token entropy. Experiments across datasets of varying sizes, different training frameworks, multiple benchmarks, and on both base and instruct models demonstrate that the advantage of GIFT over SFT is universal rather than case-specific, indicating the strong generalizability of our method.
@@ 4.3 Ablation Study
In this subsection, we conduct ablation studies to validate the design choices of GIFT, focusing on three aspects: the necessity of using entropy as the metric, the validity of the theoretically derived loss function, and the empirical necessity of applying the square root of entropy during training.
4.3.1 Necessity of Using Entropy as the Metric
@@
+ GIFT 7.8 (+2.3) 24.6 (+8.6) 78.0 (+1.3) 33.0 (+0.6) 35.8 (+3.2)
Table 2: Comparison between entropy-based importance weighting and NLL-based importance weighting. As shown in the table, the SFT algorithm with NLL-based importance weighting performs significantly worse than the entropy-based variant, and on many tasks even underperforms the original model.
To demonstrate the necessity of entropy-based importance weighting, we compared our method with a baseline using the negative log-likelihood (NLL). Unlike entropy-based importance weighting, NLL emphasizes tokens that appear less frequently rather than those that are difficult to predict or carry more information. Assuming the logits are $z_{i}$, the masking rate $\beta_{i}$ for the $i$-th token under the NLL setting is:
$\beta_{i}=-\log\bigl(\text{softmax}(z_{i})^{x_{0}^{i}}\bigr)$ (13)
@@
4.3.2 Validity of the Theoretically Derived Loss Function
@@
$L=-\sum_{i}\mathbb{E}_{t}\left[\mathbf{1}\!\left[x_{t}^{i}=\mathbf{M}\right]\frac{w_{i}}{t}\log p\left(x_{0}^{i}\mid x_{t}\right)\right]$ (14)
@@
+ GIFT (SW) 5.3 (-0.2) 23.8 (+7.8) 75.9 (-0.8) 31.6 (-0.8) 34.2 (+1.6)
+ GIFT 7.8 (+2.3) 24.6 (+8.6) 78.0 (+1.3) 33.0 (+0.6) 35.8 (+3.2)
Table 3: Comparison between GIFT, GIFT (SW) (simple token-wise weighting using our square-root entropy as the weight), and Dream. Dream yields negligible improvements. GIFT (SW) provides some gains, but still performs worse than SFT, whereas GIFT significantly outperforms SFT.
Table 3 presents a comparison between GIFT, GIFT (SW) (i.e., simple token-wise weighting using our square-root entropy as the weight), and Dream. Dream yields only minimal improvements and, in some cases, performs worse than standard SFT. GIFT (SW) provides some gains but still performs worse than SFT, whereas GIFT significantly outperforms SFT. This demonstrates that the diffusion loss derived from theoretical principles possesses superior properties.
4.3.3 Empirical Necessity of Applying the Square Root of Entropy During Training
Below, we explain why we choose the square root of entropy rather than entropy as the metric. As shown in Table 4, we observed that directly using token weights based on raw entropy leads to large variations across tokens, which can significantly reduce training stability and even cause oscillations. To address this issue, we implement GIFT using the square root of entropy as the weighting factor. This modification effectively reduces the scale of the gradient norms, resulting in more stable training and improved overall convergence behavior.
@@ 4.4 Time Efficiency Analysis
In the time efficiency analysis, there are two counteracting factors. First, GIFT requires an additional forward pass compared to SFT, resulting in a slightly higher computational overhead. However, it is important to note that the cost of a single gradient-free forward pass is significantly lower than that of a backward pass; thus, this factor only causes a marginal decrease in speed. Second, the expected number of masked tokens per iteration may differ between GIFT and SFT, which also impacts the overall training speed.
Dataset Framework SFT GIFT
s1K wd1 46 min 57 min
Tulu3 dllm 28 min 26 min
Table 5: Time efficiency comparison between SFT and GIFT on different tasks and frameworks.
As shown in Table 5, we compare the training speed of GIFT and SFT under two distinct settings: (1) LoRA fine-tuning on the s1K dataset for 20 epochs using the wd1 training framework on 4 H100 GPUs; and (2) full fine-tuning on the Tulu3-10k dataset for 5 epochs using the dllm training framework on 8 H100 GPUs. These scenarios represent representative results selected from our main experiments. The results indicate that while GIFT is merely slightly slower than SFT during LoRA fine-tuning, it is marginally faster than SFT during full fine-tuning. We attribute this to the fact that the additional forward pass incurs only a minor overhead. Conversely, since GIFT prioritizes more significant tokens, the expected total number of masked tokens may be lower than that of SFT, thereby reducing the computational load during backpropagation. This reduction in computation has a more pronounced impact during full fine-tuning compared to LoRA, leading to the observation that our method achieves faster speeds than SFT in the full fine-tuning setting.
@@ 5.1 Diffusion Large Language Models
Diffusion models, initially developed for continuous data, have recently been adapted to discrete sequences, including natural language. Early work such as Structured Denoising Diffusion Models (D3PMs) (Austin et al., 2021) extended discrete diffusion to text, showing competitive performance in character-level generation and large vocabulary language modeling. DiffusionBERT (He et al., 2023) combined masked language modeling with diffusion, improving generative quality over conventional masked LMs. Later, masked diffusion frameworks (Sahoo et al., 2024; Shi et al., 2024) introduced simplified and generalized training objectives that reduce complexity while maintaining performance, and support state-dependent masking strategies. Diffusion Large Language Models (dLLMs) including LLaDA (Nie et al., 2025) and Dream  (Ye et al., 2025) demonstrated that diffusion-based models can scale to billions of parameters, achieving strong context learning and instruction-following capabilities. Techniques such as variance-reduced preference optimization (Zhu et al., 2025) and reinforcement of lateral thought chains (Huang et al., 2025) further enhance generative stability and reasoning ability, while efficient perplexity bounds and ratio matching (Haxholli et al., 2025) improve sampling efficiency. More recently, works such as d1 (Zhao et al., 2025) and wd1 (Tang et al., 2025) have explored the possibility of applying reinforcement learning on top of existing dLLMs. LLaDA 2.0 (Bie et al., 2025) pushed the boundaries of model capacity by scaling diffusion language models to 100B parameters, offering a comprehensive analysis of scaling laws.
@@ 5.2 Weighted Methods in Language Model Learning
A complementary line of research focuses on improving large language models through weighted supervised fine-tuning (SFT) and reinforcement learning. Weighted SFT techniques aim to prioritize informative tokens. DFT (Wu et al., 2025) addresses the limited generalization of SFT by dynamically rescaling token-level objectives to stabilize gradient updates. Rho-1 (Lin et al., 2024) challenges the conventional uniform next-token prediction objective by selectively training on useful tokens identified via a reference model. Similarly, Luo et al., 2023 propose a re-weighting strategy for active learning in NER, assigning dynamic smoothed weights to tokens to boost performance. In reinforcement learning, applying weighting to improve learning effectiveness is also a common strategy. Cui et al., 2025 investigate the entropy dynamics of reinforcement learning for reasoning LLMs, and propose entropy-based weighting methods to prevent entropy collapse and encourage exploration. Wang et al., 2025 highlight the critical role of high-entropy minority tokens in RLVR, demonstrating that restricting updates to these tokens significantly enhances reasoning capabilities. Together, these approaches demonstrate that token-level weighting is an important and effective way to improve the optimization of language models.
@@ 6 Conclusion
We introduced GIFT, a theoretically grounded importance-aware fine-tuning algorithm for diffusion language models. By leveraging token-level entropy as a measure of importance, GIFT prioritizes high-uncertainty tokens and preserves diffusion process properties. Experiments across varying datasets, different training frameworks, multiple benchmarks, and on both base and instruct models demonstrate the universal advantage of the GIFT method over SFT. These results highlight entropy-based importance weighting as a simple yet powerful strategy for enhancing training effectiveness and stability in diffusion language models.
Limitations
We identify two primary limitations of this work. First, our theoretical derivation relies on certain necessary simplifying assumptions. Specifically, we assume a fixed matrix $Q$ during the derivation; however, in practical entropy-driven masking, $Q$ functions more as a heuristic variance-control scheme. We encourage future research to extend the theoretical framework to account for a dynamic $Q$. Second, the scale of our experimental validation is still limited. This is primarily due to computational constraints. Our computational resources restricted us from using even larger models, more extensive datasets, or a higher number of experimental trials.
Ethical Considerations
Potential Risks.
This paper aims to advance the field of machine learning. The work is solely centered on the methodology itself; how it is applied and for what purpose is entirely up to the users. As with other large-model research, models trained with our algorithm may generate content that is factually incorrect or potentially harmful, and such outputs should be carefully vetted by users.
Usage of Large Language Models.
In this paper, LLMs were used for polishing and improving phrases and sentences, but they did not participate in any aspects of idea conception, experimental design, theoretical proofs, figure creation, or paper structure design.
@@ Appendix A Proof of the Weighted SFT Loss
In order to prove this theorem, we will first prove two lemmas, and then proceed to prove the main theorem. Our proof follows that of Ou et al., 2025, with the key difference that our $Q$ matrix incorporates varying masking rates $\beta$, whereas Ou et al., 2025 assumes all masking rates are identical.
Here, we assume a fixed matrix $Q$ during the derivation; however, in practical entropy-driven masking, $Q$ functions more as a heuristic variance-control scheme. We encourage future research to extend the theoretical framework to account for a dynamic $Q$.
@@ A.1 Proof of Lemmas
$\frac{p_{t}(x_{t}^{1},\dots,{x^{\prime}}^{i}_{t},\dots,x_{t}^{d})}{p_{t}(x_{t}^{1},\dots,x^{i}_{t},\dots,x_{t}^{d})}$
$\quad=\frac{\exp(-\beta_{x_{t}^{i}}\,\bar{f}(t))}{1-\exp(-\beta_{x_{t}^{i}}\,\bar{f}(t))}\;p_{0}({x^{\prime}}_{t}^{i}\mid x_{t}^{\mathrm{unmasked}}).$ (17)
@@
$\frac{p_{t}(x_{t}^{1},\dots,{x^{\prime}}^{i}_{t},\dots,x_{t}^{d})}{p_{t}(x_{t}^{1},\dots,x^{i}_{t},\dots,x_{t}^{d})}=0;$ (18)
@@
$=\prod_{i=0}^{n-1}\Big(1+Q_{s_{i}}(x_{s_{i}},x_{s_{i}})(s_{i+1}-s_{i})$
$\quad+o(s_{i+1}-s_{i})\Big)$
$=\exp\Big(\sum_{i=0}^{n-1}\ln\Big(1+Q_{s_{i}}(x_{s_{i}},x_{s_{i}})$
$\quad\times(s_{i+1}-s_{i})+o(s_{i+1}-s_{i})\Big)\Big)$
$=\exp\Big(\sum_{i=0}^{n-1}[f(s_{i})Q(x_{s_{i}},x_{s_{i}})(s_{i+1}-s_{i})$
$\quad+o(s_{i+1}-s_{i})]\Big)$
Now we select a series of appropriate $n$ and $s_{1},s_{2},\dots,s_{n-1}$ such that while $n\rightarrow+\infty$ , $\max(s_{i+1}-s_{i})\rightarrow 0$. Now by recalling the definition of integral, we obtain:
$\exp\Big(\sum_{i=0}^{n-1}[f(s_{i})Q(x_{s_{i}},x_{s_{i}})(s_{i+1}-s_{i})$
$\quad+o(s_{i+1}-s_{i})]\Big)\rightarrow\exp(-\beta_{x_{0}}\bar{f}(t))$
@@
$p_{t|0}(x_{t}|x_{0})=\begin{cases}\exp(-\beta_{x_{0}}\bar{f}(t))\\ \quad\text{if }x_{t}=x_{0}\\ 1-\exp(-\beta_{x_{0}}\bar{f}(t))\\ \quad\text{if }x_{t}=[M]\\ 0\\ \quad\text{otherwise}\end{cases}$
Now, based on the above result, recalling the independence of different dimensions in the diffusion process, we can get a stronger result:
@@
Suppose that we have $x_{t}$ in which the tokens $a_{1},a_{2},...,a_{k}$ are masked and $b_{1},b_{2},...,b_{n-k}$ are unmasked; then:
$p_{t}(x_{t})$ $=\prod_{i=1}^{k}(1-\exp(-\beta_{a_{i}}\bar{f}(t)))$
$\quad\times\prod_{j=1}^{n-k}\exp(-\beta_{b_{j}}\bar{f}(t))p_{0}(x_{t}^{\text{unmasked}})$ (19)
@@
$p_{t}([M]\dots[M]x_{t}^{k+1}\dots x_{t}^{n})$
$=\sum_{x_{0}}p_{t|0}([M]\dots[M]x_{t}^{k+1}\dots x_{t}^{n}|x_{0})p_{0}(x_{0})$
$=\sum_{x_{0}^{1}\dots x_{0}^{n}}\prod_{i=1}^{k}p^{k}_{t|0}([M]|x_{0}^{k})$
$\quad\times\prod_{i=k+1}^{n}p_{t|0}(x_{t}^{k}|x_{0}^{k})p_{0}(x_{0}^{1},\dots x_{0}^{n})$
$=\sum_{x_{0}^{1}\dots x_{0}^{n}}\prod_{i=1}^{k}(1-\exp(-\beta_{i}\bar{f}(t)))$
$\quad\times\prod_{i=k+1}^{n}\exp(-\beta_{i}\bar{f}(t))p_{0}(x_{0}^{1},\dots x_{0}^{n})$
$=\sum_{x_{0}^{k+1}\dots x_{0}^{n}}\prod_{i=1}^{k}(1-\exp(-\beta_{i}\bar{f}(t)))$
$\quad\times\prod_{i=k+1}^{n}\exp(-\beta_{i}\bar{f}(t))p_{0}(x_{0}^{k+1},\dots x_{0}^{n})$
Recalling the defination, $p_{0}(x_{0}^{k+1},\dots,x_{0}^{n})$ can be expressed as $p_{0}(x_{t}^{\text{unmasked}}),$ the desired result follows.